i have to find an equation of the line with a slope of -4/3 and a y-intrercept 3?
i have to leave it in slope-intercept form.

Answer by mathslover(132)

The general equation of a line in the slope intercept form is
given by
y= mx + c where m is the slope and c is the y-intercept
in the given question slope = -4/3 and y-intercept is 3. replacing these values in the general equation we have,
y= -4/3 x + 3